# GraphLoRA: Structure-Aware Contrastive Low-Rank Adaptation for Cross-Graph Transfer Learning

### v1

@@ Abstract.
Graph Neural Networks (GNNs) have demonstrated remarkable proficiency in handling a range of graph analytical tasks across various domains, such as e-commerce and social networks. Despite their versatility, GNNs face significant challenges in transferability, limiting their utility in real-world applications. Existing research in GNN transfer learning overlooks discrepancies in distribution among various graph datasets, facing challenges when transferring across different distributions. How to effectively adopt a well-trained GNN to new graphs with varying feature and structural distributions remains an under-explored problem. Taking inspiration from the success of Low-Rank Adaptation (LoRA) in adapting large language models to various domains, we propose GraphLoRA, an effective and parameter-efficient method for transferring well-trained GNNs to diverse graph domains. Specifically, we first propose a Structure-aware Maximum Mean Discrepancy (SMMD) to align divergent node feature distributions across source and target graphs. Moreover, we introduce low-rank adaptation by injecting a small trainable GNN alongside the pre-trained one, effectively bridging structural distribution gaps while mitigating the catastrophic forgetting. Additionally, a structure-aware regularization objective is proposed to enhance the adaptability of the pre-trained GNN to target graph with scarce supervision labels. Extensive experiments on six real-world datasets demonstrate the effectiveness of GraphLoRA against eleven baselines by tuning only $20\%$ of parameters, even across disparate graph domains. The code is available at https://anonymous.4open.science/r/GraphLoRA.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Figure 1. Negative transfer occurs in cross-graph adaption, where PubMed, CiteSeer, and Cora are citation networks, whereas Photo and Computer are co-purchase networks.
Despite their superiority in capturing intricate graph relationships, GNNs face significant challenges when confronting substantial variations in underlying data distributions (Cao et al., 2023; Liu et al., 2022). Transferring a well-trained GNN to another graph typically yields suboptimal performance (i.e., negative transfer (Zhao et al., 2024)) due to out-of-distribution issues. As depicted in Figure 1, negative transfer occurs in cross-graph adaption, even when transferring between citation networks or co-purchase networks within a same domain.
@@ 4.2. Node Feature Adaptation
where $\boldsymbol{T}$ represents the transition matrix, which is related to the adjacency matrix $\boldsymbol{A}^{t}$, and $\psi_{r}$ represents the weight coefficient. We utilize a popular variant of the diffusion matrix, Personalized PageRank (PPR) (Page et al., 1998), which employs $\boldsymbol{T}=\boldsymbol{A}^{t}\boldsymbol{D}^{-1}$ and $\psi_{r}=\alpha\left(1-\alpha\right)^{r}$, where $\boldsymbol{D}$ denotes the diagonal degree matrix, i.e. $\boldsymbol{D}_{i,i}=\sum_{j=1}^{N^{t}}{\boldsymbol{A}_{i,j}^{t}}$, and $\alpha\in\left(0,1\right)$ represents the teleport probability. The elements $\boldsymbol{S}_{i,j}$ in the obtained diffusion matrix $\boldsymbol{S}$ indicate the strength of relationships between node $v_{i}^{t}$ and node $v_{j}^{t}$.
@@ 4.5. Optimization
GRACEt-PM 79.44±0.15 75.80±0.16 67.74±0.05 57.40±0.00 76.82±0.24 64.44±0.30 92.14±0.14 85.95±0.10 84.81±0.41 76.15±0.09
4.6. Complexity Analysis
5. Experiments
  • RQ1:
    How effective and efficient is GraphLoRA?
  • RQ2:
    Is GraphLoRA sensitive to hyperparameters?
  • RQ3:
    How do different modules contribute to its effectiveness?
  • RQ4:
    Can GraphLoRA mitigate catastrophic forgetting?
  • RQ5:
    Can GraphLoRA learn more distinguishable node representations than other baselines?
5.1. Experimental Setup
5.1.1. Datasets
We evaluate GraphLoRA on six publicly available datasets: three citation networks (PubMed, CiteSeer, and Cora) (Yang et al., 2016), two co-purchase networks (Photo and Computer) (Shchur et al., 2018), and one large-scale network (Reddit) (Hamilton et al., 2017). The statistics and detailed descriptions of these datasets are provided in Appendix B.
5.1.2. Baselines
We compare GraphLoRA with various baselines, including supervised methods GCN (Kipf and Welling, 2017) and GAT (Velickovic et al., 2018), graph contrastive learning methods GRACE (Zhu et al., 2020), COSTA (Zhang et al., 2022b), CCA-SSG (Zhang et al., 2021), and HomoGCL (Li et al., 2023b), as well as graph prompt learning method GPPT (Sun et al., 2022), and transfer learning methods GRACEt, GTOT (Zhang et al., 2022a), AdapterGNN (Li et al., 2024) and GraphControl (Zhu et al., 2023). Among these, GRACEt involves pretraining a GNN on the source graph using GRACE and then transferring it to the target graph for testing. Detailed descriptions of the baselines are provided in Appendix C.
5.1.3. Settings
For GraphLoRA, we use a standard two-layer GAT model as the backbone. The projector $p\left(\cdot;\boldsymbol{\omega}\right)$ and classifier $c(\cdot)$ are implemented with a single linear layer. We pre-train the GNN using GRACE and fine-tune it on the target graph using our method. Experiments are conducted in the public setting with sufficient labels, and in the 5-shot and 10-shot settings with limited labels. In the public setting, we partition citation networks using public splits (Yang et al., 2016), where each category has only 20 available labels for training. For co-purchase networks and Reddit, we randomly split the dataset into training (10%), validation (10%), and testing (80%) sets. In the 5-shot and 10-shot settings, each category in the training set contains only 5 and 10 labels, respectively. For both scenarios, 80% of the dataset is allocated for testing, while the remaining data is used as the validation set. For all methods, we conduct the experiments five times and report the average classification accuracy and standard deviation. The settings for baselines and additional results are provided in the Appendix due to space constraints.
5.2. Performance Comparison (RQ 1)
The performance of GraphLoRA on node classification tasks in both the public and 10-shot settings is presented in Table 1. Due to space constraints, we present only the experimental results of transfer learning methods on the pre-training datasets PubMed and Photo. The comprehensive table can be found in Appendix E.1, Table 7. GraphLoRA achieves either the best or second-best performance in most cases, underscoring its effectiveness. Compared to non-transfer learning scenarios, transfer learning scenarios are more challenging. Nevertheless, GraphLoRA achieves an average improvement of 1.01% over the best baseline results and 3.33% over GRACE. Specifically, it achieves an average improvement of 2.23% over GRACE in the public setting and 4.43% in the 10-shot setting. GraphLoRA shows a more significant performance improvement in the 10-shot setting, underscoring its effectiveness in scenarios with scarce labels.
5.2.1. Cross-graph Transfer Learning
(a) Average accuracy across different shots.
(b) GraphLoRA improvements over GRACE and GRACEt.
Figure 3. Experimental results across different shots.
Method PubMed CiteSeer Cora Photo Computer Avg.
GCN 2.1s 1.7s 2.3s 2.3s 2.1s 2.1s 11.6s 11.2s 12.6s 12.1s 6.0s
GAT 2.8s 2.8s 2.8s 9.0s 2.5s 8.0s 14.3s 5.5s 16.1s 4.4s 6.8s
CCA-SSG 8.7s 5.9s 6.8s 7.8s 4.5s 5.0s 4.9s 6.9s 13.4s 9.6s 7.3s
GRACE 88.5s 88.7s 6.4s 19.2s 16.9s 18.8s 31.7s 24.8s 64.6s 62.2s 42.2s
GRACEt 0.6s 0.4s 11.5s 12.0s 0.6s 12.0s 12.1s 4.1s 12.9s 5.8s 7.2s
GTOT 17.8s 79.3s 9.8s 17.6s 6.2s 10.9s 23.2s 19.9s 79.3s 56.3s 32.0s
GraphControl 0.9s 1.1s 0.3s 1.5s 1.0s 0.8s 15.6s 9.8s 29.5s 13.0s 7.4s
5.2.2. Scarce Labeling Impact on Performance
5.3. Efficiency Comparison (RQ 1)
Table 3. Experimental results on the Reddit dataset.
Method PM→R CS→R C→R P→R C→R R→R
Variants PubMed CiteSeer Cora Photo Computer
Figure 4. Performance across varying hyperparameter values.
5.4. Results on Large-Scale Dataset (RQ 1)
GraphLoRA can be easily applied to large-scale graphs using sampling techniques. We compare it with other transfer learning methods on the large-scale Reddit dataset, using the same sampling technique for all methods. The results are presented in Table 3, indicating that GraphLoRA performs well on large-scale graphs. Specifically, GraphLoRA outperforms all other methods in transfer learning scenarios and is second only to AdapterGNN in the non-transfer learning scenario. Although AdapterGNN excels in non-transfer learning scenarios, it performs poorly in transfer learning scenarios. Overall, GraphLoRA demonstrates superior performance, highlighting its effectiveness on large-scale graphs.
5.5. Hyperparameter Analysis (RQ 2)
5.5.1. Impact of $\lambda$
5.5.2. Impact of $r$
The hyperparameter $r$ determines the parameter size of GraphLoRA. We evaluate GraphLoRA’s performance across $r$ values ranging from $2^{0}$ to $2^{8}$ in the public setting. The experimental results are depicted in Figure 4. The figure reveals that GraphLoRA maintains stable performance across different values of $r$. Even when $r$ is set to 1, GraphLoRA exhibits commendable performance, aligning with the understanding that a small $r$ value is adequate for LoRA (Hu et al., 2022). Notably, GraphLoRA experiences a decline in performance when $r$ is too small or too large. Generally, optimal performance is achieved when $r$ falls within the range of $2^{3}$ to $2^{5}$, with the tunable parameter ranging from 7% to 20%. This can be attributed to a small $r$ limiting parameters for effective fine-tuning, whereas a large $r$ may lead to overfitting due to an abundance of tunable parameters.
5.6. Ablation Studies (RQ 3)
To evaluate the effectiveness of each module in GraphLoRA, we compare it with seven different model variants. Specifically, ”w/ mmd” represents the method using the target term $\mathcal{L}_{mmd}$ rather than $\mathcal{L}_{smmd}$. ”w/o smmd”, ”w/o cl”, and ”w/o str” represent methods without using the target terms $\mathcal{L}_{smmd}$, $\mathcal{L}_{cl}$, and $\mathcal{L}_{str}$, respectively. Additionally, ”w/o lrd” is the method without employing low-rank decomposition for weight updates, while ”w/o nfa” and ”w/o sktl” represent methods without utilizing the node feature adaptation module and structural knowledge transfer learning module, respectively. The experimental results conducted in the public setting and 10-shot setting, following pre-training on the PubMed dataset, are depicted in Table 4. The results in the 5-shot setting can be found in Appendix E.3.
5.7. Catastrophic Forgetting Analysis (RQ 4)
[…]
FT 79.52±0.16 71.80±1.32 75.34±1.34 64.90±1.54 59.58±2.61
GTOT 76.48±1.12 78.28±0.33 72.28±0.94 67.82±2.50 61.58±2.07
GraphLoRA 80.86±0.39 79.84±0.28 79.82±0.24 80.06±0.28 79.88±0.28
(a) GRACE
(b) GRACEt
(c) GraphLoRA
Figure 5. Visualization of node embeddings on CiteSeer.
5.8. Visualization of Representations (RQ 5)
6. Conclusion
In this paper, we investigated the challenging problem of cross-graph transfer in graph neural networks. Drawing inspiration from the successful application of LoRA in fine-tuning large language models for diverse tasks and domains, we proposed a cross-graph transfer learning framework, GraphLoRA, for effective and parameter-efficient fine-tuning of pre-trained GNNs. We addressed the discrepancies in node feature distribution and graph structure between the source and target graphs by incorporating the node feature adaptation and structural knowledge transfer learning modules. Furthermore, we proposed a structure-aware regularization objective to enhance the adaptability of pre-trained GNNs for scenarios with scarce labels. Theoretical analysis guarantees that the GraphLoRA method possesses powerful representation capabilities and can precisely fit any target GNN under mild conditions. Extensive experiments on six real-world datasets demonstrate the effectiveness of GraphLoRA, even across disparate graph domains.
Appendix A Theoretical Analysis
A.1. Notations
We define $[N]\coloneqq\{1,2,\dots,N\}$. The SVD decomposition of matrix $\boldsymbol{W}$ is given as $\boldsymbol{W}=\boldsymbol{UDV}^{T}$, and $\sigma_{i}(\boldsymbol{W})=D_{i,i}$. The best rank-$r$ approximation (in the Frobenius norm or the 2-norm) of $\boldsymbol{W}$ is $\sum_{i=1}^{r}\sigma_{i}(\boldsymbol{W})\boldsymbol{u}_{i}\boldsymbol{v}_{i}^{T}$, where $\boldsymbol{u}_{i}$ and $\boldsymbol{v}_{i}$ are the $i$-th column of $\boldsymbol{U}$ and $\boldsymbol{V}$, respectively (Eckart and Young, 1936; MIRSKY, 1960; Zeng and Lee, 2023). Following (Zeng and Lee, 2023), we define the best rank-$r$ approximation as $LR_{r}(\boldsymbol{W})$. When $r\geqslant rank(\boldsymbol{W})$, it is obvious that $LR_{r}(\boldsymbol{W})=\boldsymbol{W}$. Considering the differences in architectures among various GNNs, for the sake of analytical simplicity, we consider the following general GNN architecture:
where $\boldsymbol{B}^{l}=\boldsymbol{b}^{l}\mathbf{1},\mathbf{1}=[1,1,\dots,1]\in\mathbb{R}^{1\times N}$, and $\boldsymbol{P}$ represents the message transformation matrix associated with the adjacency matrix. For analytical simplicity, we assume that $(\boldsymbol{W}^{l})_{l=1}^{L}\in\mathbb{R}^{D\times D}$ and $(\boldsymbol{b}^{l})_{l=1}^{L}\in\mathbb{R}^{D\times 1}$.
where $\overline{L}\leqslant L$.
We define an $L$-layer width-$D$ fully connected neural network (FNN) as follows: $FNN_{L,D}(\cdot;(\boldsymbol{W}_{l})_{l=1}^{L},(\boldsymbol{b}_{l})_{l=1}^{L})\coloneqq\\ \mathrm{ReLU}(\boldsymbol{W}_{L}\mathrm{ReLU}(\boldsymbol{W}_{L-1}\mathrm{ReLU}(\ldots)+\boldsymbol{b}_{L-1})+\boldsymbol{b}_{L})$, where $(\boldsymbol{W}_{l})_{l=1}^{L}\in\mathbb{R}^{D\times D}$ are weight matrices, and $(\boldsymbol{b}_{l})_{l=1}^{L}\in\mathbb{R}^{D}$ are bias vectors. The target FNN $\overline{f}$, frozen FNN $f_{0}$, and adapted FNN $f$ are defined as follows:
where $\overline{L}\leqslant L$.
A.2. Expressive Power of Fully Connected Neural Networks with LoRA
Assumption 1.
Lemma 0.
Lemma 0.
Lemma 0.
The detailed proofs of the aforementioned lemmas can be found in (Zeng and Lee, 2023), specifically within the proofs of Lemma 3, Theorem 3, and Theorem 5.
A.3. Expressive Power of Graph Neural Networks with LoRA
Assumption 2.
Theorem 4.
Proof.
First, let’s revisit the expression of the GNN: $\boldsymbol{H}^{l}=\mathrm{ReLU}(\boldsymbol{W}^{l}\boldsymbol{H}^{l-1}\boldsymbol{P}+\boldsymbol{B}^{l})$. We can observe that the GNN expression is akin to the FNN expression. However, there are two differences between the expressions of GNN and FNN. First, the input of the FNN is a single sample, while the input of the GNN includes all samples. Second, the GNN requires message passing at each layer: $\bar{\boldsymbol{H}}^{l}=\boldsymbol{H}^{l-1}\boldsymbol{P}$. We set $\boldsymbol{W}_{l}=\boldsymbol{W}^{l}$ for $l\in[L]$ and $\overline{\boldsymbol{W}}_{i}=\overline{\boldsymbol{W}}^{i}$ for $i\in[\overline{L}]$. Under Assumption 2, it is straightforward to observe that $(\boldsymbol{W}_{l})_{l=1}^{L}$ and $(\overline{\boldsymbol{W}}_{i})_{i=1}^{\overline{L}}$ satisfy Assumption 1. Despite the aforementioned differences between FNN and GNN, the proof process of Lemma 2 (Zeng and Lee, 2023) demonstrates that these differences do not affect the proof of the conclusion. Consequently, we can easily deduce that Theorem 4 holds. ∎
Theorem 5.
(24) $\mathbb{E}\|g(\boldsymbol{X})-\overline{g}(\boldsymbol{X})\|_{2}\leqslant\xi^{{}^{\prime}}\sum_{i=1}^{\overline{L}}{\max_{k\in[\overline{L}]}}(\|\overline{\boldsymbol{W}}^{k}\|_{2}+E^{k})^{\overline{L}-i}E^{i}\|\boldsymbol{P}\|_{2}^{\overline{L}-i+1}.$
Proof.
Similarly, we set $\boldsymbol{W}_{l}=\boldsymbol{W}^{l}$ for $l\in[L]$ and $\overline{\boldsymbol{W}}_{i}=\overline{\boldsymbol{W}}^{i}$ for $i\in[\overline{L}]$. Under Assumption 2, it is straightforward to observe that $(\boldsymbol{W}_{l})_{l=1}^{L}$ and $(\overline{\boldsymbol{W}}_{i})_{i=1}^{\overline{L}}$ satisfy Assumption 1. Following the proof of Lemma 3 (Zeng and Lee, 2023) and substituting the expression of GNN for the expression of FNN in the proof process of Lemma 3 (Zeng and Lee, 2023), it is easy to deduce that Theorem 5 holds. ∎
Appendix B Datasets
B.1. Dataset Statistics
Table 6. Statistics of datasets.
Dataset #Nodes #Edges #Features #Classes
PubMed 19,717 88,651 500 3
CiteSeer 3,327 9,228 3,703 6
Cora 2,708 10,556 1,433 7
Photo 7,650 238,163 745 8
Computer 13,752 491,722 767 10
Reddit 232,965 114,615,892 602 41
B.2. Dataset Descriptions
  • Citation networks. PubMed, CiteSeer, and Cora are three standard citation network benchmark datasets. In these datasets, the nodes correspond to research papers, and the edges represent the citations between papers. The node features are derived from the bag-of-words representation of the papers, while the node labels indicate the academic topics of the papers.
  • Co-purchase networks. The Photo and Computer datasets are segments of the Amazon co-purchase graph (McAuley et al., 2015). In these datasets, the nodes represent products, while the edges represent the frequent co-purchasing relationship between two products. The node features are derived from the bag-of-words representation of product reviews, and the labels indicate the categories of the products.
Method PubMed CiteSeer Cora Photo Computer
GCN 78.66±0.56 73.28±0.93 70.50±0.75 64.52±1.40 82.00±0.97 71.88±1.26 92.17±0.75 85.39±1.57 86.66±1.20 71.97±1.16
GRACEt-PM 79.44±0.15 75.80±0.16 67.74±0.05 57.40±0.00 76.82±0.24 64.44±0.30 92.14±0.14 85.95±0.10 84.81±0.41 76.15±0.09

### v2

@@ Abstract.
Graph Neural Networks (GNNs) have demonstrated remarkable proficiency in handling a range of graph analytical tasks across various domains, such as e-commerce and social networks. Despite their versatility, GNNs face significant challenges in transferability, limiting their utility in real-world applications. Existing research in GNN transfer learning overlooks discrepancies in distribution among various graph datasets, facing challenges when transferring across different distributions. How to effectively adopt a well-trained GNN to new graphs with varying feature and structural distributions remains an under-explored problem. Taking inspiration from the success of Low-Rank Adaptation (LoRA) in adapting large language models to various domains, we propose GraphLoRA, an effective and parameter-efficient method for transferring well-trained GNNs to diverse graph domains. Specifically, we first propose a Structure-aware Maximum Mean Discrepancy (SMMD) to align divergent node feature distributions across source and target graphs. Moreover, we introduce low-rank adaptation by injecting a small trainable GNN alongside the pre-trained one, effectively bridging structural distribution gaps while mitigating the catastrophic forgetting. Additionally, a structure-aware regularization objective is proposed to enhance the adaptability of the pre-trained GNN to target graph with scarce supervision labels. Extensive experiments on eight real-world datasets demonstrate the effectiveness of GraphLoRA against fourteen baselines by tuning only $20\%$ of parameters, even across disparate graph domains. The code is available at https://github.com/AllminerLab/GraphLoRA.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709186isbn: 979-8-4007-1245-6/25/08ccs: Computing methodologies Learning latent representationsccs: Information systems Data mining
@@ 1. Introduction
Figure 1. Negative transfer occurs in cross-graph adaptation, where PubMed, CiteSeer, and Cora are citation networks, whereas Photo and Computer are co-purchase networks.
Despite their superiority in capturing intricate graph relationships, GNNs heavily rely on graph labels, which are often insufficient in real-world scenarios (Sun et al., 2023). Transfer learning is a common solution to address the issue of label sparsity (Zhuang et al., 2021). However, GNNs face significant challenges in transfer learning due to substantial variations in underlying data distributions (Cao et al., 2023; Liu et al., 2022). Transferring a well-trained GNN to another graph typically yields suboptimal performance (i.e., negative transfer (Zhao et al., 2024)) due to out-of-distribution issues. As depicted in Figure 1, negative transfer occurs in cross-graph adaptation, even when transferring between citation networks or co-purchase networks within the same domain.
@@ 4.2. Node Feature Adaptation
where $\boldsymbol{T}$ represents the transition matrix, which is related to the adjacency matrix $\boldsymbol{A}^{t}$, and $\psi_{r}$ represents the weight coefficient. We utilize a popular variant of the diffusion matrix, Personalized PageRank (PPR) (Page et al., 1998), which employs $\boldsymbol{T}=\boldsymbol{A}^{t}\boldsymbol{D}^{-1}$ and $\psi_{r}=\alpha\left(1-\alpha\right)^{r}$, where $\boldsymbol{D}$ denotes the diagonal degree matrix, i.e. $\boldsymbol{D}_{i,i}=\sum_{j=1}^{N^{t}}{\boldsymbol{A}_{i,j}^{t}}$, and $\alpha\in\left(0,1\right)$ represents the teleport probability. The elements $\boldsymbol{S}_{i,j}$ in the obtained diffusion matrix $\boldsymbol{S}$ indicate the strength of relationships between node $v_{i}^{t}$ and node $v_{j}^{t}$. For PPR, the diffusion matrix has the closed-form expression:
(7) $S=\alpha(I-(1-\alpha)D^{-1/2}AD^{-1/2})^{-1}.$
@@ 4.4. Structure-aware Regularization
GPF 79.48±0.51 74.78±1.39 71.32±0.26 66.76±0.68 82.10±0.26 76.30±0.51 91.61±0.60 86.76±1.55 77.44±1.37 70.48±1.96
GraphPrompt 75.23±0.93 74.27±1.44 69.71±1.06 65.88±0.91 79.90±0.74 75.02±0.56 86.35±0.41 84.06±0.89 72.43±0.27 66.54±0.60
ProG 75.85±0.45 71.43±0.98 71.31±0.99 68.48±1.26 82.03±0.59 76.69±0.83 85.18±1.70 86.86±0.62 66.65±1.95 67.20±1.54
transfer GRACEt-PM 79.44±0.15 75.80±0.16 67.74±0.05 57.40±0.00 76.82±0.24 64.44±0.30 92.14±0.14 85.95±0.10 84.81±0.41 76.15±0.09